OpenStudy (anonymous):

I Need Help :( x + 5y = 3 2x + 9y = 4 Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations Part 2: Provide the solution to the system

OpenStudy (dumbcow):

x = 3-5y 2(3-5y) +9y = 4 6 -10y +9y = 4 -y = -2 y = 2 x = 3-5(2) x = -7

OpenStudy (anonymous):

Karen in the first equation find value of x in terms of y ie. x= 3-5y put it in the second equation ....so u will get the equation in terms of y.

OpenStudy (anonymous):

Thanxx =]